# RAPTOR: A Foundation Policy for Quadrotor Control

### v1

@@ Acknowledgments
J.E. formulated the main ideas, implemented them, conducted the experiments and wrote the paper. D.A. and G.L. provided supervision and guided the direction during all phases of the project and helped writing the paper.
@@
The code and dataset of sampled quadrotor dynamics as well as the trained teacher checkpoints are available from https://doi.org/10.5281/zenodo.17096679  (?) and the Git repository https://github.com/rl-tools/raptor  (?). Please also refer to the project page at https://raptor.rl.tools  (?) where an interactive simulation can be used to test the RAPTOR policy. The full-length and quality video can also be accessed at https://youtu.be/hVzdWRFTX3k  (?).
@@
Corresponding author. Email: jonas.eschmann@berkeley.com
@@
$\alpha$ $=e^{\frac{\Delta t}{T}}$ (S3)
@@
We find that this mitigation significantly reduces the z-axis oscillations on the non-EKF-based platforms.
@@
We establish a physically plausible posterior distribution over quadrotors through the following relations:
@@
A simple example would be a reference trajectory with a jump in linear velocity. From the perspective of a stateful policy that has only been trained on position control (no trajectory dynamics in the observation space) this abrupt change in linear velocity appears like a quadrotor with an unrealistically high thrust-to-weight ratio, unrealistically fast angular response and/or even a world with much larger than $1$ G gravity. The latter happens if e.g., the velocity step is downwards in z and the quadrotor is not upside down or if the action inputs have been low/zero.
To counter-act this, we train the expert and student policies using a simple probabilistic mixture model of trajectories. With $50\%$ probability the task is just tracking the null-trajectory (position control, going back to the origin from any initial state). In the other $50\%$ the task is to track randomly sampled trajectories.
We would like to cover a wide variety of possible reference trajectories. Trajectories are vector-valued functions of time so we need to design a broad distribution over functions. Here we take inspiration from the Gaussian Process (GP) community, which has been concerned with designing prior distributions over functions since its inception. Additionally, we require the sampling of reference trajectories to not slow down the simulation too much. Hence we choose to sample the reference trajectories from a second-order Langevin process. A second-order Langevin process corresponds to a GP with a certain structure in the kernel (depending on the parameters of the Langevin process)  (?, ?) but we can easily sample it incrementally while simulating the quadrotor dynamics.
Based on the results described in Section 2.4.1, we find that this approach works well for moderately agile trajectory tracking and even generalizes from second-order Langevin random walks to cyclical Lissajous trajectories. But we acknowledge that, while the second-order Langevin process covers the space of smooth functions relatively well, many trajectories with real-world use-cases like e.g., step-functions/responses are not covered or exponentially unlikely (e.g., cyclical trajectories).
@@
We do not perform variational inference in a case-by-case basis but instead in an amortized fashion, where the inference itself is conducted by the recurrent neural network policy at test time. This amortization makes the inference very compute efficient and allows us to deploy the foundation policy onto even the tiniest microcontrollers while meeting real-time constraints at high frequencies.
@@
  • Context Window Extrapolation: At inference time we would like the policy to fly for longer time than the context window used during training. During training the context is usually limited to not slow down the process too much. CNNs, by design, have a limited context window that cannot be extended. Similarly attention, being a set-to-set mapping, requires position embeddings which complicate context window extrapolation.
@@
We find that a surprisingly small (in terms of number of parameters and compute requirements) three-layer recurrent neural network is sufficient to express the desired behavior described in Section 1. Figure 7B shows the architecture which contains a dense input layer, Gated Recurrent Unit (GRU) layer and a dense output layer. The initialization of the recurrent state (value at step 0) is all zeros and we feed back the previous output of the policy as an input. Due to the small hidden dimensions the foundation policy only has $2084$ parameters:
@@
We train the foundation policy using Meta-Imitation Learning where it acts as the student. During Meta-Imitation Learning, we want to adjust the students weights to minimize the KL divergence (also known as relative entropy) between the predicted optimal action distribution by the student and the optimal action distribution (approximated by the particular teacher for the current system dynamics):
@@
Hence we want to find the log probabilities of the action distributions that we approximated as Gaussians before:
@@
Since we are aiming at quadrotors, both, the numerator and denominator multivariate Gaussian are 4 dimensional and the constants cancel.
@@
Hence, under the mild assumption of unit standard deviations, we can conduct the Meta-Imitation Learning in a principled manner by minimizing the relative differential entropy between the (approximated) optimal action distribution and the predicted optimal action distribution by the student. In practice, we implement this by using Eq. Meta-Imitation Learning Objective, which is identical to the Mean-Squared Error (MSE), as the loss function and by optimizing the student policy’s parameters $\boldsymbol{\theta}$ using gradient descent.
@@
Note
All these movies are combined in https://youtu.be/hVzdWRFTX3k

### v2

@@ Acknowledgments
J.E. formulated the main ideas, implemented them, conducted the experiments and wrote the paper. D.A. and G.L. provided supervision and guided the direction during all phases of the project and helped write the paper.
@@
The code and dataset of sampled quadrotor dynamics as well as the trained teacher checkpoints are available from https://doi.org/10.5281/zenodo.17096679 (?) and the Git repository https://github.com/rl-tools/raptor (?). Please also refer to the project page at https://raptor.rl.tools (?) where an interactive simulation can be used to test the RAPTOR policy. The full-length, high-quality video can also be accessed at https://youtu.be/hVzdWRFTX3k (?).
@@
Corresponding author. Email: jonas.eschmann@berkeley.edu
@@
Quadrotor Dynamics
The L2F simulator (?) uses the following equations of motion:
$\dot{\mathbf{p}}=$ $\;\mathbf{v}$
$\dot{\mathbf{q}}=$ $\;\mathbf{q}\circ\left[0\;\;\boldsymbol{\omega}/2\right]^{\top}$
$\dot{\mathbf{v}}=$ $\;\frac{1}{m}\left(\mathbf{\mathbf{R}}(\mathbf{q})\left(\sum_{i=0}^{3}\mathbf{r}_{f_{i}}f_{i}\right)+\mathbf{f}_{\text{ext}}\right)+\mathbf{g}$
$f_{i}=$ $\;\sum_{j=0}^{2}c_{f_{j}}\left(\omega_{m_{i}}\right)^{j}$
$\dot{\boldsymbol{\omega}}=$ $\;\mathbf{J}^{-1}\left(\boldsymbol{\tau}+\left(\mathbf{J}\boldsymbol{\omega}\right)\times\boldsymbol{\omega}\right)$
$\boldsymbol{\tau}=$ $\;\sum_{i=0}^{3}\left(\mathbf{r}_{p_{i}}\times\mathbf{r}_{f_{i}}\right)f_{i}+\mathbf{r}_{\tau_{i}}c_{m}f_{i}$
$\mathbf{r}_{p_{i}}=$ $\;\left[\pm\frac{l_{\text{arm}}}{\sqrt{2}},\pm\frac{l_{\text{arm}}}{\sqrt{2}},0\right]^{\top}$
$\mathbf{r}_{\tau_{i}}=$ $\;\left[0,0,\pm 1\right]^{\top}$
[…]
$\boldsymbol{\omega}_{sp}:=$ $\;\mathbf{a}$
Where $\boldsymbol{\omega}$ is in the body frame. We only consider the case of planar quadrotors, where $\mathbf{r}_{f_{i}}=\left[0,0,1\right]^{\top}\;\forall i\in 0\ldots 3$. $\mathbf{r}_{p_{i}}$ and $\mathbf{r}_{\tau_{i}}$ are set up according to the Crazyflie motor layout conventions: (front-right, back-right, back-left, front-left) and (CCW, CW, CCW, CW), respectively. Note that this corresponds to the top-down view and the body frame is following the Front-Left-Up (FLU) conventions. CW corresponds to a torque direction of $\mathbf{r}_{\tau_{i}}=\left[0,0,1\right]$. When deploying on platforms with other motor-layout conventions, a re-mapping of the motor commands is applied.
@@
$\alpha$ $=e^{-\frac{\Delta t}{T}}$ (S3)
@@
We find that this mitigation substantially reduces the z-axis oscillations on the non-EKF-based platforms.
@@
We establish a physically plausible generative distribution over quadrotors through the following relations:
@@
Finally, we sample the standard deviation used for sampling the random disturbance force $\mathbf{f}_{\text{ext}}$ as $10\%$ of the surplus thrust $r_{\text{t2w}}-1$:
$\sigma_{\mathbf{f}_{\text{ext}}}$ $\sim\mathrm{Uniform}(0,(r_{\text{t2w}}-1)\cdot 0.1)$ (S27)
The idea behind choosing $10\%$ is that by applying the $3\sigma$-rule, the sampled random disturbance will be confined at $30\%$ of the surplus thrust with a $\sim 99.7\%$ probability. This makes it exceedingly unlikely that we sample forces that are too strong to be compensated by the control policy.
@@
A simple example would be a reference trajectory with a jump in linear velocity. From the perspective of a stateful policy that has only been trained on position control (no trajectory dynamics in the observation space) this abrupt change in linear velocity appears like a quadrotor with an unrealistically high thrust-to-weight ratio, unrealistically fast angular response and/or even a world with much larger than $1$ G gravity. The latter happens if, for example, the velocity step is downwards in z and the quadrotor is not upside down or if the action inputs have been low/zero.
To counter-act this, we train the expert and student policies using a simple probabilistic mixture model of trajectories. With $50\%$ probability, the task is just tracking the null-trajectory (position control, going back to the origin from any initial state). In the other $50\%$ the task is to track randomly sampled trajectories. Note that position and linear velocity are just offset by the trajectory ($\mathbf{p}:=\mathbf{p}-\mathbf{p}_{\text{target}}$, $\mathbf{v}:=\mathbf{v}-\mathbf{v}_{\text{target}}$) in the observations as well as in the reward function. We neither change the structure nor adjust the parameters.
We would like to cover a wide variety of possible reference trajectories. Trajectories are vector-valued functions of time, so we need to design a broad distribution over functions. Here we take inspiration from the Gaussian Process (GP) community, which has been concerned with designing prior distributions over functions since its inception. Additionally, we require the sampling of reference trajectories to not slow down the simulation too much. Hence, we choose to sample the reference trajectories from a second-order Langevin process. A second-order Langevin process corresponds to a GP with a certain structure in the kernel (depending on the parameters of the Langevin process) (?, ?) but we can easily sample it incrementally while simulating the quadrotor dynamics.
Based on the results described in the Trajectory Tracking subsection, we find that this approach works well for moderately agile trajectory tracking and even generalizes from second-order Langevin random walks to cyclical Lissajous trajectories. But we acknowledge that, although the second-order Langevin process covers the space of smooth functions relatively well, many trajectories with real-world use-cases such as step-functions/responses are not covered or exponentially unlikely (such as cyclical trajectories).
@@
We do not perform variational inference on a case-by-case basis but instead in an amortized fashion, where the inference itself is conducted by the recurrent neural network policy at test time. This amortization makes the inference very compute efficient and allows us to deploy the foundation policy onto even the tiniest microcontrollers while meeting real-time constraints at high frequencies.
@@
  2. 2.
    Context Window Extrapolation: At inference time we would like the policy to fly for longer time than the context window used during training. During training the context is usually limited to not slow down the process too much. CNNs, by design, have a limited context window that cannot be extended. Similarly, attention, being a set-to-set mapping, requires position embeddings which complicate context window extrapolation.
  3. 3.
@@
We find that a surprisingly small (in terms of number of parameters and compute requirements) three-layer recurrent neural network is sufficient to express the desired behavior described in the Introduction section. Figure 7B shows the architecture which contains a dense input layer, Gated Recurrent Unit (GRU) layer and a dense output layer. The initialization of the recurrent state (value at step 0) is all zeros and we feed back the previous output of the policy as an input. Due to the small hidden dimensions the foundation policy only has $2084$ parameters:
@@
We train the foundation policy using Meta-Imitation Learning where it acts as the student. During Meta-Imitation Learning, we want to adjust the student’s weights to minimize the KL divergence (also known as relative entropy) between the predicted optimal action distribution by the student and the optimal action distribution (approximated by the particular teacher for the current system dynamics):
@@
Hence, we want to find the log probabilities of the action distributions that we approximated as Gaussians before:
@@
Since we are aiming at quadrotors, both, the numerator and denominator multivariate Gaussian are 4-dimensional and the constants cancel.
@@
Hence, under the mild assumption of unit standard deviations, we can conduct the Meta-Imitation Learning in a principled manner by minimizing the relative differential entropy between the (approximated) optimal action distribution and the predicted optimal action distribution by the student. In practice, we implement this by using Eq. S54, which is identical to the Mean-Squared Error (MSE), as the loss function and by optimizing the student policy’s parameters $\boldsymbol{\theta}$ using gradient descent.